OpenStudy (anonymous):

If you help me, i'll give you a medal :D Find the exact value by using a half-angle identity. tan(7pi/8)

OpenStudy (jdoe0001):

$$\bf \cfrac{7\pi}{8} \times 2 \implies \cfrac{7\pi}{4}\\ \cfrac{\frac{7\pi}{4}}{2} \implies \cfrac{7\pi}{8}\\ tan\left(\cfrac{\frac{7\pi}{4}}{2}\right) = \cfrac{sin\left(\frac{7\pi}{4}\right)}{1+cos\left(\frac{7\pi}{4}\right)} $$ so, what are the values for sine and cosine for \(\bf \cfrac{7\pi}{4}\ \ ?\)

OpenStudy (anonymous):

sin is sqrt 2 and cos is - sqrt 2

OpenStudy (jdoe0001):

\(\bf \cfrac{7\pi}{4}\) is in the 4th quadrant in the 4th quadrant sine is negative and cosine is positive, otherwise the values are correct

OpenStudy (jdoe0001):

or you can use this Unit Circle instead -> http://i.stack.imgur.com/r8uHr.gif

OpenStudy (anonymous):

ooooooooooh I forgot that part :p

OpenStudy (jdoe0001):

\(\bf tan\left(\cfrac{\frac{7\pi}{4}}{2}\right) = \cfrac{sin\left(\frac{7\pi}{4}\right)}{1+cos\left(\frac{7\pi}{4}\right)}\\ tan\left(\cfrac{\frac{7\pi}{4}}{2}\right) =\large \cfrac{-\frac{\sqrt{2}}{2}}{1+\frac{\sqrt{2}}{2}}\)

OpenStudy (jdoe0001):

what would that give you?

OpenStudy (anonymous):

- sqrt 2/ 1 + 2sqrt2?

OpenStudy (jdoe0001):

yes, as you'd noticed in the Unit Circle, those are the values for cosine and sine at \(\bf \cfrac{7\pi}{4}\)

OpenStudy (anonymous):

Ok, so that would be the answer?

OpenStudy (jdoe0001):

well, what does \(\bf \large \cfrac{-\frac{\sqrt{2}}{2}}{1+\frac{\sqrt{2}}{2}}\) give you?

OpenStudy (anonymous):

I'm not sure. Do you have to change the 1?

OpenStudy (jdoe0001):

keep in mind that, in the denominator => \(\bf \large 1+\frac{\sqrt{2}}{2} \implies \frac{1}{1}+\frac{\sqrt{2}}{2}\) , so is just a plain fractions addition

OpenStudy (anonymous):

so it's \[-\sqrt{2}/ ((2+\sqrt{2})/ 2)\]

OpenStudy (jdoe0001):

well, the numerator should be negative, as in \(\bf -\cfrac{\sqrt{2}}{2}\) the denominator is fine though

OpenStudy (jdoe0001):

\(\bf \large \bf \cfrac{-\frac{\sqrt{2}}{2}} {\frac{2+\sqrt{2}}{2}} \implies -\frac{\sqrt{2}}{2} \times \frac{2}{2+\sqrt{2}} \)

OpenStudy (anonymous):

\[-(2\sqrt{2})/ 4 +2\sqrt{2}\]

OpenStudy (jdoe0001):

well \(\bf \large -\frac{\sqrt{2}}{\cancel{2}} \times \frac{\cancel{2}}{2+\sqrt{2}} \implies \cfrac{-\sqrt{2}}{2+\sqrt{2}}\)

OpenStudy (anonymous):

Thank you! :D

OpenStudy (jdoe0001):

one sec

OpenStudy (jdoe0001):

$$\bf \text{now simplifying the denominator by multiplying for the conjugate}\\ \cfrac{-\sqrt{2}}{2+\sqrt{2}} \times \cfrac{2-\sqrt{2}}{2-\sqrt{2}}\\ \text{keep in mind that}\\ (a-b)(a+b) = (a^2-b^2)\\ \cfrac{-\sqrt{2} \times (2-\sqrt{2})}{(2)^2-(\sqrt{2})^2} $$

OpenStudy (jdoe0001):

$$\bf \frac{2-2\sqrt{2}}{4-2} \implies \frac{\cancel{2}-\cancel{2}\sqrt{2}}{\cancel{2}} $$

OpenStudy (jdoe0001):

so, that leaves you with a simplified version of \(\bf 1-\sqrt{2}\)

OpenStudy (anonymous):

Oooooooooooooooooooh! Thank you!!!

OpenStudy (jdoe0001):

yw